OpenStudy (anonymous):

a car is travelling with uniform acceleration along a straight road. the road has marker posts every 100m.when the car passes one post, it has a speed of 10m/s and, when it passes the next one its speed is 20 m/s. find acceleration.

OpenStudy (anonymous):

@sauravshakya i got wrong answer help please

OpenStudy (anonymous):

Hint: v^2=u^2+2as

OpenStudy (anonymous):

v=20m/s u=10m/s and s=100m

OpenStudy (anonymous):

Now, find a

OpenStudy (anonymous):

i got 1.5

OpenStudy (anonymous):

Ya.... thats very correct

OpenStudy (anonymous):

didn't we have to do for 1st post in 1st case where u=0 n v= 10 m 2nd case as u did...and adding it will give the result

OpenStudy (anonymous):

Isnt 1.5 the answer

OpenStudy (anonymous):

1st i got 2.0 and now 1.5 ..got it just calculation mistake .

OpenStudy (anonymous):

Oh...... the thing u said cannot be done because we dont know if it was at rest at starting post or not

OpenStudy (anonymous):

yes got it ...thanx